Circular No. 9104

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2010 A1 (HILL) R. E. Hill, Lunar and Planetary Laboratory, reports his discovery of

a comet on CCD images taken with the Catalina 0.68-m Schmidt reflector (discovery observation tabulated below), describing the object from coadded exposures in 2′′–4′′ seeing as having a nuclear condensation ∼ 4′′–5′′ across with a 12′′–15′′-radius coma beyond that and no obvious tail (but a slight east-west elongation) in moonlight; his follow-up images in thin haze on Jan. 7.24–7.25 UT show a similar appearance but with a short, fanshaped tail 20′′–30′′ long in p.a. ∼ 270^{o}. Four co-added images obtained by E. C. Beshore with the Mt. Lemmon 1.5-m reflector on Jan. 6.5 showed a very faint tail perhaps 15′′ long in p.a. 270^{o}. Following posting on the Minor Planet Center’s ‘NEOCP’ webpage, other CCD astrometrists have noted the object’s cometary appearance. Y. Ikari (Moriyama, Shiga-Ken, Japan, 0.26-m f/7 reflector, Jan. 6.6) measured a coma diameter of 0′.2. P. Birtwhistle (Great Shefford, Berkshire, England, 0.40-m f/6 Schmidt-Cassegrain reflector, Jan. 7.1) found a diffuse coma of size 15′′ × 22′′, elongated in p.a. 310^{o}, with a weak condensation offset to the southeast (the brightest part of the surrounding coma is 6′′ × 12′′ and extends in p.a. 50^{o} and 230^{o}). H. Sato (Ota-ku, Tokyo, Japan, remotely using a 0.25-m f/3.4 reflector of the RAS Observatory, near Mayhill, NM, U.S.A., Jan. 7.2) finds a 20′′ coma elongated toward p.a. 110^{o}. W. H. Ryan (Magdalena Ridge Observatory, 2.4-m f/8.9 reflector, Jan. 7.3) remarked on a distinct coma.

2010 UT α_{2000} δ_{2000} Mag. Observer
Jan. 6.46381 9^{h}11^{m}26.^{s}44 +33^{o}08′34′′.6 17.6 Hill

Further astrometry, the following preliminary parabolic orbital elements, and an ephemeris are given on MPEC 2010-A31. It is possible that the comet is of short period.

T = 2010 June 24.055 TT  ω = 219.^{o}047
Ω = 6.140 }2000.0
q = 1.02031 AU  i = 11.782

COMET 81P/WILD Visual total-magnitude estimates: Nov. 24.17 UT, 10.5 (J. J. Gonzalez,

Leon, Spain, 20-cm reflector); Dec. 12.16, 10.3 (Gonzalez); 29.21, 10.3 (M. Meyer, Limburg, Germany, 15-cm refractor).

2010 January 7  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2010 CBAT  Daniel W. E. Green